Firm will profit 400k at the end year 6. G= 11 %  Discount rate = 18%, you will attain half of firm. What amount willing to invest?
Answer #1

value of firm today=(profit at the end of year 6/(r-g))/(1+r)^5

=(400/(18%-11%))/(1+18%)^5

=2497.77K OR 2497766.95

What amount willing to invest for 50% share=2497766.95/2=1248883.48 or 1248.88K

the above is answer..